How long will it take $8000 to triple itself if it is invested at 6% simple interest.


Answer by Edwin McCravy(6939)


How long will it take $8000 to triple itself if it is
invested at 6% simple interest.
For $8000 to triple, it must become $24000
A = P + Prt
24000 = 8000 + 8000(.06)t
24000 = 8000 + 480t
16000 = 480t
Divide both sides by 480
33.33 = t
Answer: 33 years and 4 months.
That's not keeping up with inflation because
33 years ago a gallon of gas was less than 40 cents.
Edwin